OpenStudy (anonymous):

please solve this one. thanks. 2/3d + 5/2d = 4

OpenStudy (anonymous):

d=24/19

OpenStudy (anonymous):

do you have a solution?

OpenStudy (anonymous):

that is the solution...

OpenStudy (anonymous):

okay. thank you :D

OpenStudy (anonymous):

you should understand the steps 2/3d + 5/2d = 19/6d and now you have something you can work with easier. You multiply both side by 6d and then you get 19 = 24d and then you divide both sides by 24 to get 19/24 not 24/19.

OpenStudy (anonymous):

No. (19/6)d=4 ==> 6(19/6)d=6*4 ==> 19d=24 ==> d=24/19

OpenStudy (anonymous):

wolfram alpha it dude your wrong

OpenStudy (anonymous):

You can always check your answer by plugging in your solution. So, (2/3)(24/19) + (5/2)(24/19) = 16/19 + 60/19 = 4

OpenStudy (anonymous):

d = 19/24

OpenStudy (anonymous):

Unfortunately, (2/3)(19/24) + (5/2)(19/24) = 19/36 + 95/48 = 361/144 = 2.5069

OpenStudy (anonymous):

omg you are either purposely trying to screw this person up or you are just really dumb

OpenStudy (anonymous):

(2/(3(19/24))+(5/(2(19/24))) = 4 You are not even entering the variable into the equation the right way that is why it gave the wrong answer

OpenStudy (anonymous):

The right way? It depends on how you specify the equation. I know that people who see \[(2/3)x\]on their paper and type online as \[2/3x\]which you see as \[2/(3x)\] BUT \[2\div3\times x = (2x)\div3\]

OpenStudy (anonymous):

I am tired of arguing with you. It is clear you are not understanding order of operations. 2/3*x DOES NOT EQUAL 2/(3x)

jhonyy9 (jhonyy9):

2/3 d +5/2 d =4 so than 6 will be the denominator common and will get (4d +15d)/6 = 24/6 4d+15d=24 19d=24 d=24/19

jhonyy9 (jhonyy9):

hope so much that is understandably

OpenStudy (anonymous):

it depends how you look at the question Jesscama it 2/(3d) + 5/(2d) = 4 or is it (2/3)d + (5/2)d = 4

OpenStudy (anonymous):

is it**

jhonyy9 (jhonyy9):

ok you are right but for this is necessary using parentheses so when dont us d is considered that is part of numerator